OpenStudy (anonymous):

Which answer best describes the solution to the equation −12j + 4 = −4 – 12j? A. one solution B. no solution C. an infinite number of solutions D. j = 3 I think it is A.

OpenStudy (anonymous):

j is generally used in electronics for vectors, as opposed to maths, which uses i

OpenStudy (anonymous):

-12j +4 =-4 -12j -12j +12j = -4 -4 0j = -8 No solution